# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v3

@@ 3.1.1. Recursive Construction:
where $G_{1}^{\prime}=\mathbb{F}_{3}(\alpha)$. In addition, let $C_{11}^{\prime}=\langle(1,2,1,2,1),(\alpha,\alpha^{2},\alpha^{3},\alpha^{4},\alpha^{5})\rangle$ be the $5,2,4]_{3^{5}}-$code, $C_{11}^{\prime\prime}=C_{11}^{{}^{\prime}\perp_{E}}\subset(G_{1}^{\prime\prime})^{5}$ the $5,3,3]_{3^{5}}-$code, and $C_{11}=\langle(1,1,1,0,0),(1,2,0,1,0)\rangle$ the self-orthogonal $5,2,3]_{3}-$code, which are the constituent codes of the QC code $\mathcal{C}_{1}$. Thus, we may obtain the $m^{u}\ell$-length QC codes
(54) $\mathcal{C}_{u}\cong 11^{u-1}C_{11}^{\prime}\oplus 11^{u-1}C_{11}^{{}^{\prime}\perp}\oplus\mathcal{C}_{u-1},\mbox{ for }u\geq 2.$
According to Corollary 3.4, it is possible to get a infinite family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of ESO QC codes with parameters $\left[5\cdot 11^{u},\frac{5\cdot 11^{u}-1}{2},\geq 3\cdot 11^{u-1}\right]_{3}$.
3.2. Constructing ESO/ESD/EDC QC Codes with Square-Root-Like Lower Bound
Self-dual codes have applications in several mathematical problems such as in the construction of designs, lattices, and quantum error-correcting codes (see Subsection 2.4). Notably, the construction of infinite families of $n_{i},k_{i},d_{i}]_{q}$ self-dual codes is itself an interesting problem. Additionally, code families where $d_{i}\geq c\sqrt{n_{i}}$ for some constant $c$ are said to satisfy the square-root-like lower bound. In the special case when $c=1$, we say that the code family satisfies the square-root lower bound. According to [infiniteselfdualfamilies], there are very few constructions of infinite families of self-dual codes whose minimum distances meet the square-root lower bound, which are depicted in that paper. Notably, the work in  [Calderbank] presents a unique construction of an infinite family of self-dual QC codes.
To the best of our knowledge, the authors of [infiniteselfdualfamilies] in 2025 presented the first infinite families of non-binary ESD (and HSD) cyclic codes satisfying the square-root-like lower bound to their minimum distances. Using results from [lingsole] and developments presented in this work (see Corollary 3.4), we construct several infinite families of ESD QC codes which satisfy the square-root like lower bound. Moreover, we also construct HSD QC codes using results in the next subsection. We first state the following helpful lemma.
@@ 3.2. Constructing ESO/ESD/EDC QC Codes with Square-Root-Like Lower Bound
Let $m$ and $u$ be positive integers. If $u\geq 2$, then $\sqrt{m^{u}}\leq m^{u-1}.$
Proof.
$u\geq 2\Rightarrow u+2\leq 2u\Rightarrow m^{u+2}\leq m^{2u}\Rightarrow m^{u}\leq m^{(u-1)^{2}}\Rightarrow\sqrt{m^{u}}\leq m^{u-1}$. ∎
Let $\{\mathcal{C}_{u}\}_{u\geq 1}$ be the infinite family of QC codes proposed in Corollary 3.4. As mentioned in the statement of that corollary, $\mathcal{C}_{u}$ being ESO/EDC/ESD comes from the duality of the last constituent code over $\mathbb{F}_{q}$. Based on this observation, we have the following result.
@@
Let $R_{q,m}^{\ell}\cong\bigoplus_{i=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{i}(x)\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{i}^{*}(x)\rangle}\right)^{\ell}\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$, where $(g_{i}(x),g_{i}^{*}(x))$ are pairs of distinct reciprocal polynomials, for all $1\leq i\leq t$. Let $\mathcal{C}_{1}\subset\mathbb{F}_{q}^{m\ell}$ be the ESD $\ell$-QC code with constituent codes $C_{1i}\oplus C_{1i}^{\perp}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{i}(x)\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{i}^{*}(x)\rangle}\right)^{\ell}$ and a ESD code $C_{1,t+1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$, where
$d(C_{11})\geq d^{\perp}(C_{11})\geq...\geq d(C_{1t})\geq d^{\perp}(C_{1t})\geq d(C_{1,t+1})\geq c\sqrt{\ell}.$
Then there exists an infinite family $\{\mathcal{C}_{u}\}_{u\geq 2}$ of ESD QC codes with parameters
(55) $\left[m^{u}\ell,\ell\left(\frac{m^{u}-1}{m-1}\right)\left(\sum_{j=1}^{t}\deg(g_{j})\right)+k_{1s},\geq m^{u-1}d_{1,t+1}\right],$
where all these codes have square-root-like bound on their minimum distances.
Proof.
According to Corollary 3.4 and the fact that the last constituent code of $\mathcal{C}_{1}$ is self-dual, it follows that all $\mathcal{C}_{u}$ is self-dual as well, for $u\geq 1$. Additionally, from Lemma 3.6, for each $m^{u}\ell$-length QC code $\mathcal{C}_{u}$, with $u\geq 2$, we have
(56) $d(\mathcal{C}_{u})\geq m^{u-1}d_{1,t+1}\geq\sqrt{m^{u}}d_{1,t+1}\geq\sqrt{m^{u}}c\sqrt{\ell}=c\sqrt{m^{u}\ell},$
namely, the minimum distances in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound, and the result follows. ∎
Remark 4.
@@
where $G_{1}^{\prime}=\mathbb{F}_{5}(\alpha)$. Let $C_{11}^{\prime}$ be a $6,3,4]_{5^{5}}-$GRS code, and once $C_{11}^{{}^{\prime\prime}}=C_{11}^{{}^{\prime}\perp_{E}}$, then $C_{11}^{\prime\prime}$ is also a $6,3,4]_{5^{5}}-$GRS code. In addition, let $C_{11}=\langle(1,0,0,2,2,4),(0,1,0,2,4,2),(0,0,1,4,2,2)\rangle$ be a self-dual $6,3,4]_{5}$ code. Notice all constituent codes are MDS. Therefore, according to Theorem 3.7, it is possible to get a infinite family $\{\mathcal{C}_{u}\}_{u\geq 2}$ of ESD QC codes with parameters $\left[4\cdot 11^{u},2\cdot 11^{u},\geq 4\cdot 11^{u-1}\right]$ for $u\geq 2$, with square-root-like bound for their minimum distances.
@@
We conclude this subsection by emphasizing that, alongside the seminal constructions of infinite families of self-dual cyclic codes satisfying the square-root-like bound in [infiniteselfdualfamilies], our work introduces a novel construction of infinite families of QC codes that also attain this square-root-like lower bound on minimum distance.
3.3. Galois Closed QC Codes and Their Duals
@@
Having QC codes whose Hermitian and Euclidean duals are equal is a valuable property as it allows us to move freely between the two inner products. To achieve this property, we introduce the notion of the Galois closure of a code $C$: By  [Jin2010], given a positive integer $m\geq 1$ and $c=(c_{1},c_{2},...,c_{n})\in C\subset\mathbb{F}_{q}^{n}$ (in this case, $C$ is just a subset), we define $c^{m}=(c_{1}^{m},c_{2}^{m},...,c_{n}^{m})$ and $C^{m}=\{c^{m}:c\in C\}\subset\mathbb{F}_{q}^{n}.$ If $C$ is a subspace of $\mathbb{F}_{q}^{n}$ and $m$ is a power of the characteristic of $\mathbb{F}_{q}$, i.e., $m=q^{t}$ for some $t\in\mathbb{Z}^{+}$, then $C^{m}$ is also a subspace of $\mathbb{F}_{q}^{n}$.
@@
Let $\mathbb{F}_{q}\subset\mathbb{F}_{q^{r}}$ and $C$ be an $\mathbb{F}_{q^{r}}$-code. We call $C$ Galois closed if $C=C^{q}$. See [Bierbrauer] for further information. The Frobenius map $C\mapsto C^{q}$ respects the quasi-cyclic property of $C$ so that $C^{q}$ is also quasi-cyclic whenever $C$ is. Furthermore, if $\phi(C),\phi(C^{q})\subset\left(\frac{\mathbb{F}_{q^{2}}[x]}{\langle x^{m}-1\rangle}\right)^{\ell}$, then we can look at the constituent codes of $C$ and $C^{q}$ and assume that the irreducible factors of $x^{m}-1$ are aligned for both decompositions.
Remark 5.
@@
For any code $C$ over $\mathbb{F}_{q^{2}}$, we observe that the Hermitian dual $C^{\perp_{H}}$ is equal to the Euclidean dual $\left(C^{q}\right){{}^{\perp_{E}}}$ of $C^{q}$. Hence, $C$ is HSO if and only if $C\subset(C^{q})^{\perp_{E}}$, i.e., $C^{q}\subset C^{\perp_{E}}$. Moreover, if $C$ is Galois closed, then $C^{\perp_{H}}=C^{\perp_{E}}$.
@@
The following two results give a classification and construction for the Galois closed codes in $\mathbb{F}_{q^{2}}^{\ell}$, namely that they are direct sums of one-dimensional subspaces spanned by vectors with entries 0 or $v_{i}$ satisfying $v_{i}^{q-1}=\beta$ for some fixed $\beta\in\mathbb{F}_{q^{2}}$. This allows for easy construction of Galois closed codes and demonstrates that they are not rare but rather plentiful, making them indeed useful and desirable.
@@
Proposition 3.11.
Let $C\subset\mathbb{F}_{q^{2}}$ be a Galois closed code. $C$ is HSO code if and only if $C$ is ESO.
Proof.
$C$ is HSO if and only if $C^{q}\subset C^{\perp_{E}}$. Since $C^{q}=C$, the result follows. ∎
In general, $C\neq C^{q}$, but they have the same parameters. According to Remark 5, if $C$ is Galois closed, then its Hermitian and Euclidean duals are exactly the same, i.e. $C^{\perp_{H}}=C^{\perp_{E}}$. Furthermore, the following theorems shows that if $C$ is Galois closed, then its Euclidean dual is also Galois closed.
@@
(57) $\mathbf{c_{\gamma}}^{q}$ $=$ $(\sum_{i=0}^{k-1}\gamma_{i}\mathbf{c_{i}})^{q}=\sum_{i=0}^{k-1}\gamma_{i}^{q}\mathbf{c_{i+1}}=\sum_{i=0}^{k-2}\gamma_{i}^{q}\mathbf{c_{i+1}}+\gamma_{k-1}^{q}\sum_{i=0}^{k-1}\beta_{i}\mathbf{c_{i}},\mbox{ where }\beta_{i}\in\mathbb{F}_{q^{r}},i=0,...,k-1;$
@@
(58) $\langle\mathbf{x}^{q},\mathbf{c}_{\gamma}\rangle_{E}=\sum_{i=0}^{k-1}\langle\mathbf{x}^{q},\gamma_{i}\mathbf{c_{i}}\rangle_{E}=\left(\sum_{i=0}^{k-1}\langle\mathbf{x},\gamma_{i}^{q^{r-1}}\mathbf{c_{i-1}}\rangle_{E}\right)^{q}=0$
where $\mathbf{c_{-1}}$ is written as linear combination of $\mathbf{c_{0}},...,\mathbf{c_{k-1}}$. ∎
The Trace Representation (Theorem 2.2) of $\ell$-QC codes depends on the constituent codes and $q$-cyclotomic cosets. It turns out that the the constituent codes of a QC code being Galois closed directly determines whether $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ is Galois closed.
@@
(59) $\mathcal{C}^{q}\cong\Bigg(\bigoplus_{j=1}^{t}({(C^{q})}{{}^{\prime}}_{j}\oplus{(C^{q})}{{}^{\prime\prime}}_{j})\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}(C^{q})_{i}\bigg).$
As the corresponding constituent codes of $\mathcal{C}^{q}$ and $\mathcal{C}$ are aligned, we have that $(C^{q})^{\prime}_{j}=C^{\prime q}_{j}$ and $(C^{q})^{\prime\prime}_{j}=C^{\prime\prime q}_{j}$ for all $1\leq j\leq t$, and $(C^{q})_{i}=C_{i}^{q}$ for each $1\leq i\leq s$. Thus $\mathcal{C}=\mathcal{C}^{q}$ as codes if and only if $(C^{q})_{i}=C_{i}^{q}=C_{i}$ for each $1\leq i\leq s$, and $(C^{q})^{\prime}_{j}=C^{\prime q}_{j}=C^{\prime}_{j}$ and $(C^{q})^{\prime\prime}_{j}=C^{\prime\prime q}_{j}=C^{\prime\prime}_{j}$ for all $1\leq j\leq t$. This can also be seen via the Trace Representation (14) in conjunction with properties of the trace map.

### v4

@@ 44 $\mathcal{C}_{4}\cong C_{4,1}\oplus C_{4,2}\oplus\cdots\oplus C_{4,2t+s}$
where $G_{1}^{\prime}=\mathbb{F}_{3}(\alpha)$. In addition, let $C_{1,1}^{\prime}=\langle(1,2,1,2,1),(\alpha,\alpha^{2},\alpha^{3},\alpha^{4},\alpha^{5})\rangle$ be the $5,2,4]_{3^{5}}-$code, $C_{1,1}^{\prime\prime}=C_{1,1}^{{}^{\prime}\perp_{E}}\subset(G_{1}^{\prime\prime})^{5}$ be the $5,3,3]_{3^{5}}-$code, and $C_{11}=\langle(1,1,1,0,0),(1,2,0,1,0)\rangle$ be the self-orthogonal $5,2,3]_{3}-$code. Let these codes be are the constituent codes of the QC code $\mathcal{C}_{1}$. For any $u\geq 2$, let the first two constituent codes of $\mathcal{C}_{u}$ be multiples of $\mathcal{C}_{1,1}^{\prime}$ and $C_{1,1}^{{}^{\prime}\perp_{E}}$ to obtain the $m^{u}\ell$-length QC codes
$\mathcal{C}_{u}\cong 11^{u-1}C_{1,1}^{\prime}\oplus 11^{u-1}C_{1,1}^{{}^{\prime}\perp_{E}}\oplus\mathcal{C}_{u-1}.$ (43)
According to Corollary 3.26, the $\{\mathcal{C}_{u}\}_{u\geq 1}$ is an infinite family of ESO QC codes with parameters
$\left[5\cdot 11^{u},\frac{5\cdot 11^{u}-1}{2},\geq 3\cdot 11^{u-1}\right]_{3}.$
4 Satisfying the Square-Root-Like Lower Bound
Using results from lingsole and developments presented in this paper, we show how the construction of codes in Section 3 satisfy the square-root-like lower bound and then present explicit examples. We first state the following helpful lemma.
@@ 4 Satisfying the Square-Root-Like Lower Bound
If $m\geq 1$ and $u\geq 2$ are positive integers, then $\sqrt{m^{u}}\leq m^{u-1}.$
Proof 4.30.
If $u\geq 2$ , then $u\leq 2u-2$. For $m\geq 1$, this gives $m^{u}\leq m^{2u-2}$ and so $\sqrt{m^{u}}\leq m^{u-1}$. $\qed$
@@
Let $\{\mathcal{C}_{u}\}_{u\geq 1}$ be the infinite family of ESO $q$-ary QC codes proposed in Theorem 3.24. Then
  • (i)
    If $d(C_{1,s})\geq c\sqrt{\ell}$ for some constant $c$, then all codes in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound for their minimum distances.
  • (ii)
    Otherwise, there exist a positive integer $L>1$ so that the all codes in the subfamily $\{\mathcal{C}_{u}\}_{u\geq L}$ of $\{\mathcal{C}_{u}\}_{u\geq 1}$ satisfies the square-root-like bound for their minimum distances.
Proof 4.32.
  • (i)
    By Lemma 4.29, for each $m^{u}\ell$-length QC code $\mathcal{C}_{u}$ with $u\geq 2$, we have
    $d(\mathcal{C}_{u})\geq m^{u-1}d_{1,s}\geq\sqrt{m^{u}}d_{1,s}\geq\sqrt{m^{u}}c\sqrt{\ell}=c\sqrt{m^{u}\ell},$
    namely, the minimum distances in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound and the result follows.
  • (ii)
    Notice that
    $m^{u-1}d_{1s}\geq c\sqrt{m^{u}\ell}\iff m^{2u-2}d_{1s}^{2}\geq c^{2}m^{u}\ell\iff m^{u-2}\geq\frac{c^{2}\ell}{d_{1s}^{2}}.$
    Hence, $u-2\geq\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}$, and when $u\geq\left\lceil\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}\right\rceil+2=:L$, the subfamily $\{\mathcal{C}_{u}\}_{u\geq L}$ of $\{\mathcal{C}_{u}\}_{u\geq 1}$ satisfies the square-root-like bound. $\qed$
In the special case where $R_{q,m}^{\ell}$ in (25) has only x1x-1 as the self-reciprocal factor of $x^{m}-1$, we obtain EDC and ESD codes that satisfying the square-root-like bound.
et $\{\mathcal{C}_{u}\}_{u\geq 1}$ be the infinite family of QC codes proposed in Corollary 3.26. As mentioned in the statement of that corollary, $\mathcal{C}_{u}$ being ESO, EDC, or ESD comes from the duality of the last constituent code over $\mathbb{F}_{q}$. Based on this observation, we have the following theorem.
Corollary 4.33.
Let $\{\mathcal{C}_{u}\}_{u\geq 1}$ be the infinite family of ESO/EDC/ESD $q$-ary QC codes proposed in Corollary 3.26. Then
  • (i)
    If $d(C_{1,t+1})\geq c\sqrt{\ell}$ for some constant $c$, then all codes in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound for their minimum distances.
  • (ii)
    Otherwise, there exist a positive integer $L>1$ so that the all codes in the subfamily $\{\mathcal{C}_{u}\}_{u\geq L}$ of $\{\mathcal{C}_{u}\}_{u\geq 1}$ satisfies the square-root-like bound for their minimum distances.
lightgray
Proof 4.34.
  • (i)
    According to Corollary 3.26 and the fact that the last constituent code of $\mathcal{C}_{1}$ is ESO (resp. EDC or ESD), it follows that all $\mathcal{C}_{u}$ are ESO (resp. EDC or ESD) as well, for $u\geq 2$. Additionally, from Lemma 4.29, for each $m^{u}\ell$-length QC code $\mathcal{C}_{u}$, with $u\geq 2$, we have
    $d(\mathcal{C}_{u})\geq m^{u-1}d_{1,t+1}\geq\sqrt{m^{u}}d_{1,t+1}\geq\sqrt{m^{u}}c\sqrt{\ell}=c\sqrt{m^{u}\ell},$
    namely, the minimum distances in $\{\mathcal{C}_{u}\}_{u\geq 2}$ satisfy the square-root-like bound and the result follows.
  • (ii)
    Notice that
    $m^{u-1}d_{1s}\geq c\sqrt{m^{u}\ell}\iff m^{2u-2}d_{1s}^{2}\geq c^{2}m^{u}\ell\iff m^{u-2}\geq\frac{c^{2}\ell}{d_{1s}^{2}}.$
    Hence, $u-2\geq\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}$, and when $u\geq\left\lceil\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}\right\rceil+2=:L$, the subfamily $\{\mathcal{C}_{u}\}_{u\geq L}$ of $\{\mathcal{C}_{u}\}_{u\geq 1}$ satisfies the square-root-like bound. $\qed$
Remark 4.35.
@@
Remark 4.36.
Additionally, is straightforward to see that the lower-bound for the minimum distance in Theorem 3.24 is asymptotically better than the square-root-like lower bound. Indeed, given that $d_{1s},m,c$, and $\ell$ are positive integers, we have
$\lim_{u\rightarrow+\infty}\frac{m^{u-1}d_{1s}}{c\sqrt{m^{u}\ell}}=\lim_{u\rightarrow+\infty}\frac{m^{\frac{u}{2}-1}d_{1s}}{c\sqrt{\ell}}=+\infty>1.$
@@
where $G_{1}^{\prime}=\mathbb{F}_{5}(\alpha)$. Let $C_{1,1}^{\prime}$ be a $6,3,4]_{5^{5}}-$GRS code and $C_{1,1}^{{}^{\prime\prime}}=C_{1,1}^{{}^{\prime}\perp_{E}}$, then $C_{11}^{\prime\prime}$ is also a $6,3,4]_{5^{5}}-$GRS code. In addition, let $C_{1,1}=\langle(1,0,0,2,2,4),(0,1,0,2,4,2),(0,0,1,4,2,2)\rangle$ be a self-dual $6,3,4]_{5}$ code. Observe that all constituent codes are MDS. Therefore, according to Corollary 4.33, it is possible to get a infinite family $\{\mathcal{C}_{u}\}_{u\geq 2}$ of ESD QC codes with parameters $\left[4\cdot 11^{u},2\cdot 11^{u},\geq 4\cdot 11^{u-1}\right]_{5}$ for $u\geq 2$ that satisfy the square-root-like bound for their minimum distances.
Example 4.38.
Let $\alpha$ be a primitive $7$-th root of unity and let
$R_{2,7}^{8}$ $\cong$ $\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{3}+x+1\rangle}\right)^{8}\oplus\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{3}+x^{2}+1\rangle}\right)^{8}\oplus\left(\frac{\mathbb{F}_{2}[x]}{\langle x-1\rangle}\right)^{8}$
Let $C_{1,1}^{\prime}$ be an $8,3,6]_{8}$ MDS code (whose construction can be checked in table), $C_{1,1}^{{}^{\prime\prime}}=C_{1,1}^{{}^{\prime}\perp_{E}}$ be an $8,5,4]_{8}$ (also MDS) code, and $C_{1,1}$ be an EDC $8,6,2]_{2}-$code with basis
$B=\{$ $(1,0,0,1,0,0,0,0),(0,1,0,1,0,0,0,0),(0,0,1,1,0,0,0,0),(0,0,0,0,1,0,0,1),(0,0,0,0,0,1,0,1),$
$(0,0,0,0,0,0,1,1)\}.$
ote that $d(D_{3})=7$, $d(D_{2,3})=3$, and $d(D_{1,2,3})=1$. So from Lemma 2 and Theorem 4, the code $\mathcal{C}_{1}$ is a $56,30,\geq 6]_{2}$ 88-QC code. The current best-known $56,30]_{2}$ code has minimum distance $10$. By Corollary 2.16, we obtain a quantum code $56,4,\geq 6]]_{2}$the best known minimum distance is $13$.
AA: I feel that the phrases I colored in red directly above short-sell our results because our lower-bound for the minimum distance is far away from the best known. G.: I agree
By Corollary 3.26, we obtain a $\left[8\cdot 7^{u},4\cdot 7^{u}+2,\geq 2\cdot 7^{u-1}\right]_{2}$ QC code. When $u=3$, $d\geq\sqrt{8\cdot 7^{u}}$, which satisfies the square-root bound.
G. Gustavo´s suggestion Note that $d(D_{3})=7$, $d(D_{2,3})=3$, and $d(D_{1,2,3})=1$. So from Lemma 2 and Theorem 4, the code $\mathcal{C}_{1}$ is a $56,30,\geq 6]_{2}$ 88-QC code. By Corollary 3.26, we obtain a $\left[8\cdot 7^{u},4\cdot 7^{u}+2,\geq 2\cdot 7^{u-1}\right]_{2}$ QC code.
When $u\geq 3$, $d\geq\sqrt{8\cdot 7^{u}}$, which shows that the infinite family $\{C_{u}\}_{u\geq 3}$ satisfies the square-root bound for the minimum distances. Moreover for $c=5$, we have $\left\lceil\log_{m}\frac{c^{2}\ell}{d_{1s}^{2}}\right\rceil+2=5$. Thus by Corollary 4.33, we then obtain an infinite family of QC codes $\{\mathcal{C}_{u}\}_{u\geq 5}$ whose minimum distances are greater than $c\sqrt{8\cdot 7^{u}}$.
@@
We conclude this section by emphasizing that, alongside the seminal constructions of infinite families of self-dual cyclic codes satisfying the square-root-like bound in infiniteselfdualfamilies, our work introduces a novel construction of infinite families of QC codes that also attain this square-root-like lower bound on minimum distance.
5 Galois Closed QC Codes and Their Duals
@@
Having QC codes whose Hermitian and Euclidean duals coincide is a valuable property, as it allows us to move freely between the two inner products. To achieve this property, we introduce the notion of the Galois closure of a code $C$ using notation from Jin2010: Given a positive integer $m\geq 1$ and $c=(c_{1},c_{2},...,c_{n})\in C\subset\mathbb{F}_{q}^{n}$ (in this case, $C$ is just a subset), we define $c^{m}=(c_{1}^{m},c_{2}^{m},...,c_{n}^{m})$ and $C^{m}=\{c^{m}:c\in C\}\subset\mathbb{F}_{q}^{n}.$ If $C$ is a subspace of $\mathbb{F}_{q}^{n}$ and $m$ is a power of the characteristic of $\mathbb{F}_{q}$, i.e., $m=q^{t}$ for some $t\in\mathbb{Z}^{+}$, then $C^{m}$ is also a subspace of $\mathbb{F}_{q}^{n}$.
@@
Let $\mathbb{F}_{q}\subset\mathbb{F}_{q^{r}}$ and $C$ be an $\mathbb{F}_{q^{r}}$-code. We say that $C$ is Galois closed if $C=C^{q}$, where further on such codes details can be found in Bierbrauer. The Frobenius map $C\mapsto C^{q}$ respects the quasi-cyclic property of $C$ so that $C^{q}$ is also quasi-cyclic whenever $C$ is. Furthermore, if $\phi(C),\phi(C^{q})\subset\left(\frac{\mathbb{F}_{q^{2}}[x]}{\langle x^{m}-1\rangle}\right)^{\ell}$, then we can look at the constituent codes of $C$ and $C^{q}$ and assume that the irreducible factors of $x^{m}-1$ are aligned for both decompositions.
Remark 5.39.
@@
For any code $C$ over $\mathbb{F}_{q^{2}}$, we observe that the Hermitian dual $C^{\perp_{H}}$ is equal to the Euclidean dual $\left(C^{q}\right){{}^{\perp_{E}}}$ of $C^{q}$. Hence, $C$ is HSO if and only if $C\subset(C^{q})^{\perp_{E}}$, i.e., $C^{q}\subset C^{\perp_{E}}$. Moreover, if $C$ is Galois closed, then $C^{\perp_{H}}=C^{\perp_{E}}$. Thus, $C$ is an HSO (resp. HDC, HSD) code if and only if $C$ is ESO (resp. EDC, ESD).
@@
The following two propositions give a classification and construction for the Galois closed codes in $\mathbb{F}_{q^{2}}^{\ell}$, namely that they are direct sums of one-dimensional subspaces spanned by vectors with entries 0 or $v_{i}$ satisfying $v_{i}^{q-1}=\beta$, for some fixed $\beta\in\mathbb{F}_{q^{2}}$. This allows for easy construction of Galois closed codes and demonstrates that they are not rare but rather plentiful, making them indeed useful and desirable.
@@
proposition Let $C\subset\mathbb{F}_{q^{2}}^{\ell}$ be a Galois closed code. Then $C$ is an HSO code if and only if $C$ is ESO.
Proof 5.42.
$C$ is HSO if and only if $C^{q}\subset C^{\perp_{E}}$. Since $C^{q}=C$, the result follows.
In general $C\neq C^{q}$, but they have the same parameters. According to Remark 5.39, if $C$ is Galois closed, then its Hermitian and Euclidean duals are exactly the same, i.e. $C^{\perp_{H}}=C^{\perp_{E}}$. Furthermore, the following theorem shows that if $C$ is Galois closed, then its Euclidean dual is also Galois closed.
Theorem 5.43.
A code $C$ over $\mathbb{F}_{q^{2}}$ is Galois closed if and only if its Euclidean dual is also Galois closed.
Proof 5.44.
We only prove the forward direction as the reverse direction is nearly identical. Let $x\in C$ and $y\in C^{\perp_{E}}$. Then
$\langle y^{q},x\rangle_{E}=\sum_{i}y_{i}^{q}x_{i}=\bigg(\sum_{i}y_{i}^{q}x_{i}\bigg)^{q}=\sum_{i}y_{i}x_{i}^{q}=\langle y,x^{q}\rangle_{E}.$
As $y\in C^{\perp_{E}}$, $y^{q}$ is in $(C^{\perp_{E}})^{q}$. Likewise, since $x\in C$ and $C=C^{q}$, we have that $\langle y,x^{q}\rangle_{E}=0$ and so $\langle y^{q},x\rangle_{E}=0$. Thus $y^{q}\in C^{\perp_{E}}$ , i.e., $y\in(C^{\perp_{E}})^{q}$, and so $C^{\perp_{E}}\subseteq(C^{\perp_{E}})^{q}$. The reverse direction follows similarly. $\qed$
@@
$\mathbf{c_{\gamma}}^{q}$ $=$ $(\sum_{i=0}^{k-1}\gamma_{i}\mathbf{c_{i}})^{q}=\sum_{i=0}^{k-1}\gamma_{i}^{q}\mathbf{c_{i+1}}=\sum_{i=0}^{k-2}\gamma_{i}^{q}\mathbf{c_{i+1}}+\gamma_{k-1}^{q}\sum_{i=0}^{k-1}\beta_{i}\mathbf{c_{i}},\mbox{ where }\beta_{i}\in\mathbb{F}_{q^{r}},i=0,...,k-1;$ (44)
@@
$\langle\mathbf{x}^{q},\mathbf{c}_{\gamma}\rangle_{E}=\sum_{i=0}^{k-1}\langle\mathbf{x}^{q},\gamma_{i}\mathbf{c_{i}}\rangle_{E}=\left(\sum_{i=0}^{k-1}\langle\mathbf{x},\gamma_{i}^{q^{r-1}}\mathbf{c_{i-1}}\rangle_{E}\right)^{q}=0$ (45)
where $\mathbf{c_{-1}}$ is written as linear combination of $\mathbf{c_{0}},...,\mathbf{c_{k-1}}$. $\qed$
The Trace Representation (lingsole, Theorem 5.1) of $\ell$-QC codes depends on the constituent codes and $q$-cyclotomic cosets. It turns out that the the constituent codes of a QC code being Galois closed directly determines whether $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ is Galois closed.
@@
$\mathcal{C}^{q}\cong\Bigg(\bigoplus_{j=1}^{t}({(C^{q})}{{}^{\prime}}_{j}\oplus{{(C^{q})}^{\prime}_{j}}^{\perp_{E}})\Bigg)\oplus\bigg(\bigoplus_{i=1}^{s}(C^{q})_{i}\bigg).$ (46)
As the corresponding constituent codes of $\mathcal{C}^{q}$ and $\mathcal{C}$ are aligned, we have that $(C^{q})^{\prime}_{j}=C^{\prime q}_{j}$ and ${(C^{q})^{\prime}_{j}}^{\perp_{E}}=((C^{\prime}_{j})^{\perp_{E}})^{q}$ for all $1\leq j\leq t$, and $(C^{q})_{i}=C_{i}^{q}$ for each $1\leq i\leq s$. Thus $\mathcal{C}=\mathcal{C}^{q}$ as codes if and only if $(C^{q})_{i}=C_{i}^{q}=C_{i}$ for each $1\leq i\leq s$, and $(C^{q})^{\prime}_{j}=C^{\prime q}_{j}=C^{\prime}_{j}$ and ${(C^{q})^{\prime}_{j}}^{\perp_{E}}=({C^{\prime}_{j}}^{\perp_{E}})^{q}={C^{\prime}_{j}}^{\perp_{E}}$ for all $1\leq j\leq t$. This can also be seen via the Trace Representation (lingsole, Theorem 5.1) in conjunction with properties of the trace map. $\qed$
We use these results to obtain the following family of $q^{2}$-ary QC codes that satisfy the square-root-like bound for their minimum distances.